# CAGE: Continuity-Aware edGE Network Unlocks Robust Floorplan Reconstruction

### v1

@@ 5.2 Comparisons to State-of-the-Art Methods
Table 1: Quantitative comparison on the Structured3D [16] test set. The running time is averaged over the whole test set. PD denotes PolyDiffuse [38] post-optimization.
Method Venue t (s) Room Corner Angle
@@
HEAT [13] CVPR22 0.11 96.9 94.0 95.4 81.7 83.2 82.5 77.6 79.0 78.3
RoomFormer [12] CVPR23 0.01 97.9 96.7 97.3 89.2 85.3 87.2 83.0 79.6 81.3
SLIBO-Net [14] NeurIPS23 0.17 99.1 97.8 98.4 88.9 82.1 85.4 87.8 81.2 84.4
PolyRoom [37] ECCV24 0.02 98.9 97.7 98.3 94.6 86.1 90.2 89.3 81.4 85.2
FRI-Net [15] ECCV24 0.09 99.5 98.7 99.1 90.8 84.9 87.8 89.6 84.3 86.9
PolyGraph [36] TVCG25 0.04 95.7 97.9 96.7 92.4 82.2 88.3 89.2 79.4 85.4
CAGE(Our) 0.01 99.6 98.7 99.1 95.0 88.6 91.7 92.5 86.4 89.3
RoomFormer+PD [38] NeurIPS23 98.7 98.1 98.4 92.8 89.3 91.0 90.8 87.4 89.1
FRI-Net+PD [15] ECCV24 99.6 98.6 99.1 94.2 88.2 91.1 91.9 86.7 89.2
Ours+PD 99.7 98.7 99.1 95.1 89.9 91.7 91.6 87.5 89.2
@@
After applying PolyDiffuse for post-optimization, our method further consolidates its leading position across all three evaluation levels, outperforming both RoomFormer and FRI-Net. Interestingly, the post-optimization with PolyDiffuse brings only marginal improvements to our model, indicating that CAGE already captures the scene structure effectively without reliance on heavy post-processing. In contrast, PolyDiffuse, as a diffusion-based model, requires substantial additional training time. Furthermore, despite not explicitly optimizing for corner predictions, our method maintains competitive performance in Corner Recall. Regarding efficiency, CAGE achieves fasted inference speed with Roomformer (both 0.01 second per image), benefiting from its lightweight architecture without the need for additional refinement stages. Overall, these results demonstrate that CAGE strikes an excellent balance between reconstruction accuracy and computational efficiency.
@@
In cross-dataset evaluation (training on Structured3D [16] and testing on SceneCAD [17]), PolyRoom and CAGE achieve the top results across Room IoU and fine-level corner/angle metrics (Table 3). PolyRoom’s segmentation-based pipeline aligns well with SceneCAD’s single-room layouts, yielding strong room-level performance, but its multi-stage design struggles with the complex multi-room structures in Structured3D (Table 1). In contrast, CAGE’s fully end-to-end, edge-based modeling consistently maintains high accuracy (85.6% Room IoU, 87.5% Corner Recall, 70.6% Angle Recall, 61.6% Angle F1) without post-processing. These results validate that combining edge-based representations with global shape modeling and denoising strategies enables CAGE to achieve robust and balanced floorplan reconstruction under domain shifts and varying scene complexities.
Table 2: Quantitative comparison on the SceneCAD validation set [17]. Results for prior methods are reported from [12, 15, 38].
@@
5.3 Ablation Study
We evaluate the impact of three core components: the edge-based polygon representation, the denoising (DN) training strategy, and the image backbone. As shown in Table 4 in Supplementary Material B.5, both the Edge and DN modules individually improve performance, with their combination yielding substantial gains—particularly on Corner and Angle metrics, with F1 improvements of +5.8% and +7.7% over the baseline. The Edge module provides the most consistent benefit, enhancing structural regularity and reducing sensitivity to missing corners. Backbone comparisons reveal that SwinTransformer-V2, when used in isolation, slightly degrades performance. However, when combined with the Edge and DN modules, it improves the Corner and Angle metrics, while ResNet-50 still achieves the highest Room F1. For a balanced trade-off across all metrics, we adopt SwinTransformer-V2 with both modules as the default configuration. Further analysis and visual examples are provided in Supplementary Material B.5.
6 Conclusion
We present CAGE, an end-to-end floorplan reconstruction framework that models rooms as sequences of edges using a dual-query transformer architecture. By predicting edges instead of corners, our method captures global geometric structure and directional priors, enabling robust inference of room layouts even from partially observed wall segments through edge-based attention. This significantly improves resilience to occlusion and layout noise. A denoising-based training strategy with perturbed and latent queries facilitates stable and efficient iterative refinement. Extensive experiments on Structured3D and SceneCAD demonstrate state-of-the-art performance. Our method still puzzles overs reconstructing scenes with dense occlusions or fine structural elements. Future work includes integrating semantic priors, extending to full architectural elements, and scaling to outdoor scenarios.
Supplementary Materials
Appendix A Implementation Details
Model Settings. Our primary image backbone is Swin Transformer V2 [45], while ResNet-50 [46] and Swin Transformer V1 [47] are used for comparison in the ablation study. Following RoomFormer [12], for ResNet-50, we extract multi-scale feature maps from the last three stages without using a feature pyramid network (FPN). For Swin Transformers, we adopt the official large-variant configurations: SwinV1 uses Swin-L pretrained weights, and SwinV2 uses SwinV2-L pretrained weights, both with a base dimension of $C=192$ and stage depths of [2, 2, 18, 2]. The window sizes differ, with SwinV1 using a window size of 12 and SwinV2 using 16. For both ResNet-50 and Swin Transformer backbones, the fourth-scale feature map is obtained via a $3\times 3$ convolution with stride 2 applied to the final-stage output. All feature maps are projected to 256 channels before being fed into the transformer. Pretrained weights from [46, 47, 45] are used for all backbone layers and are fine-tuned during training. Our transformer architecture comprises 6 encoder layers and 6 decoder layers, each with a hidden dimension of 256. We set the number of room feature codes $m$ to 20 and the number of edge queries $n$ to 40.
Appendix B Additional Results and Visualizations
B.1 Evolution of Edge Queries
To further examine the decoding process, Figure 9 presents a layer-wise visualization of edge query refinement. As decoding progresses, the predicted edge structures become increasingly coherent and better aligned with the underlying point density map, even in regions with limited or missing wall evidence. Correspondingly, the reconstructed polygons evolve from coarse outlines to precise, watertight layouts. This illustrates the effectiveness of our progressive decoding strategy in capturing both geometric accuracy and topological consistency.
B.2 More Reconstruction result on Structured3D
B.3 Failure cases on Structure3D
B.4 More Reconstruction result on SceneCAD
B.5 Extended Ablation Study
@@
We perform a comprehensive ablation study using RoomFormer [12] as the baseline to evaluate the impact of key architectural components. Specifically, we examine the contributions of two core modules—edge-based polygon representation and a denoising (DN) training strategy—as well as the influence of different image backbones. All models, including RoomFormer, are trained under identical settings for fair comparison. As summarized in Table 4 and visualized in Fig. 13, both the Edge and DN modules yield consistent improvements individually, and their combination results in further performance gains. By contrast, incorporating SwinTransformer-V2 as the backbone provides only marginal benefits beyond the combined Edge and DN configuration.
@@
To further improve learning stability and geometric reasoning, we incorporate a denoising strategy in which ground-truth data with noise are introduced alongside random queries during training. This approach supports convergence in the early stages of prediction and stablize the Hungry Matching process. As illustrated in the fourth row of Table 4, the addition of the DN module alone leads to moderate gains in Corner and Angle metrics, though Room F1 slightly drops from 97.3% to 97.1%. While the DN module is less impactful than the Edge module in isolation, its value becomes more evident in combination.
@@
We evaluate the impact of different image backbones—ResNet-50 [46], SwinTransformer-V1 [47], and SwinTransformer-V2 [45]—while keeping the proposed Edge and DN modules fixed. All configurations utilize four image scales and a token dimension of 256. As shown in the last three rows of Table 4, integrating Swin Transformer backbones yields mixed results. While SwinTransformer-V2 achieves the highest Angle F1 score (88.6%) and slightly improves Corner F1 (91.6%), these gains are relatively modest compared to the improvements brought by the Edge and DN modules. Notably, ResNet-50 combined with both modules attains the highest Room F1 score (99.2%), underscoring its effectiveness in capturing global structural context.

### v2

@@ 5.2 Comparisons to State-of-the-Art Methods
Table 1: Quantitative comparison on the Structured3D [16] test set. The running time is averaged over the whole test set. PD denotes PolyDiffuse [38] post-optimization. *: Params are calculated from the officially released models.
Method Venue Backbone Params(M)* t (s) Room Corner Angle
@@
HEAT [13] CVPR22 ResNet-50 48.9 0.11 96.9 94.0 95.4 81.7 83.2 82.5 77.6 79.0 78.3
RoomFormer [12] CVPR23 ResNet-50 40.6 0.01 97.9 96.7 97.3 89.2 85.3 87.2 83.0 79.6 81.3
SLIBO-Net [14] NeurIPS23 ResNet-50 0.17 99.1 97.8 98.4 88.9 82.1 85.4 87.8 81.2 84.4
PolyRoom [37] ECCV24 ResNet-50 42.2 0.02 98.9 97.7 98.3 94.6 86.1 90.2 89.3 81.4 85.2
FRI-Net [15] ECCV24 ResNet-50 58.6 0.09 99.5 98.7 99.1 90.8 84.9 87.8 89.6 84.3 86.9
PolyGraph [36] TVCG25 ResNet-50 0.04 95.7 97.9 96.7 92.4 82.2 88.3 89.2 79.4 85.4
CAGE(Our) SwinV2-L 211.3 0.01 99.6 98.7 99.1 95.0 88.6 91.7 92.5 86.4 89.3
RoomFormer+PD [38] NeurIPS23 ResNet-50 98.7 98.1 98.4 92.8 89.3 91.0 90.8 87.4 89.1
FRI-Net+PD [15] ECCV24 ResNet-50 99.6 98.6 99.1 94.2 88.2 91.1 91.9 86.7 89.2
Ours+PD SwinV2-L 99.7 98.7 99.1 95.1 89.9 91.7 91.6 87.5 89.2
@@
After applying PolyDiffuse for post-optimization, our method further consolidates its leading position across all three evaluation levels, outperforming both RoomFormer and FRI-Net. Interestingly, the post-optimization with PolyDiffuse brings only marginal improvements to our model, indicating that CAGE already captures the scene structure effectively without reliance on heavy post-processing. In contrast, PolyDiffuse, as a diffusion-based model, requires substantial additional training time. Furthermore, despite not explicitly optimizing for corner predictions, our method maintains competitive performance in Corner Recall. Regarding efficiency, CAGE achieves the fastest inference speed, along with RoomFormer (both 0.01 second per image), benefiting from its lightweight architecture without the need for additional refinement stages. Overall, these results demonstrate that CAGE strikes an excellent balance between reconstruction accuracy and computational efficiency.
@@
In cross-dataset evaluation (training on Structured3D [16] and testing on SceneCAD [17]), PolyRoom and CAGE achieve the top results across Room IoU and fine-level corner/angle metrics (Table 3). PolyRoom’s segmentation-based pipeline aligns well with SceneCAD’s single-room layouts, yielding strong room-level performance, but its multi-stage design struggles with the complex multi-room structures in Structured3D (Table 1). In contrast, CAGE’s end-to-end trainable, edge-based modeling consistently maintains high accuracy (85.6% Room IoU, 87.5% Corner Recall, 70.6% Angle Recall, 61.6% Angle F1) without post-processing. These results validate that combining edge-based representations with global shape modeling and denoising strategies enables CAGE to achieve robust and balanced floorplan reconstruction under domain shifts and varying scene complexities.
Table 2: Quantitative comparison on the SceneCAD validation set [17]. Results for prior methods are reported from [12; 15; 38].
@@
5.3 Ablation Study
We evaluate the impact of three core components: the edge-based polygon representation, the denoising (DN) training strategy, and the image backbone. As shown in Table 4, both the Edge and DN modules yield consistent improvements individually, and their combination results in further performance gains. By contrast, incorporating SwinTransformer-V2 as the backbone provides only marginal benefits beyond the combined Edge and DN configuration. Further analysis and visual examples are provided in Supplementary Material B.5.
@@ 5.3 Ablation Study
To further improve learning stability and geometric reasoning, we incorporate a denoising strategy in which ground-truth data with noise are introduced alongside random queries during training. This approach supports convergence in the early stages of prediction and stablize the Hungarian Matching process. As illustrated in the third row of Table 4, the addition of the DN module alone leads to moderate gains. While the DN module is less impactful than the Edge module in isolation, its value becomes more evident in combination.
@@
We evaluate the impact of different image backbones—ResNet-50 [46], SwinTransformer-V1 [47], and SwinTransformer-V2 [45]—while keeping the proposed Edge and DN modules fixed. All configurations utilize four image scales and a token dimension of 256. As shown in the last three rows of Table 4, integrating Swin Transformer backbones yields mixed results. While SwinTransformer-V2 achieves the highest Angle F1 score (89.3%) and slightly improves Corner F1 (91.7%), these gains are relatively modest compared to the improvements brought by the Edge and DN modules. Notably, ResNet-50 combined with both modules attains the highest Room F1 score (99.2%), underscoring its effectiveness in capturing global structural context.
@@
6 Conclusion
We present CAGE, an end-to-end floorplan reconstruction framework that models rooms as sequences of edges using a dual-query transformer architecture. By predicting edges instead of corners, our method captures global geometric structure and directional priors, enabling robust inference of room layouts even from partially observed wall segments through edge-based attention. This significantly improves resilience to occlusion and layout noise. A denoising-based training strategy with perturbed and latent queries facilitates stable and efficient iterative refinement. Extensive experiments on Structured3D and SceneCAD demonstrate state-of-the-art performance. Our method still puzzles over reconstructing scenes with dense occlusions or fine structural elements. Future work includes integrating semantic priors, extending to full architectural elements, and scaling to outdoor scenarios.
Supplementary Materials
Appendix A Implementation Details
Model Settings. Our primary image backbone is Swin Transformer V2 [45], while ResNet-50 [46] and Swin Transformer V1 [47] are used for comparison in the ablation study. Following RoomFormer [12], for ResNet-50, we extract multi-scale feature maps from the last three stages without using a feature pyramid network (FPN). For Swin Transformers, we adopt the official large-variant configurations: SwinV1 uses Swin-L pretrained weights, and SwinV2 uses SwinV2-L pretrained weights, both with a base dimension of $C=192$ and stage depths of [2, 2, 18, 2]. The window sizes differ, with SwinV1 using a window size of 12 and SwinV2 using 16. For both ResNet-50 and Swin Transformer backbones, the fourth-scale feature map is obtained via a $3\times 3$ convolution with stride 2 applied to the final-stage output. All feature maps are projected to 256 channels before being fed into the transformer. Pretrained weights from [46; 47; 45] are used for all backbone layers and are fine-tuned during training. Our transformer architecture comprises 6 encoder layers and 6 decoder layers, each with a hidden dimension of 256. We set the number of room feature codes $m$ to 20 and the number of edge queries $n$ to 40.
Appendix B Additional Results and Visualizations
B.1 Evolution of Edge Queries
To further examine the decoding process, we visulize a layer-wise edge query refinement as illustrated in Figure 9. As decoding progresses, the predicted edge structures become increasingly coherent and better aligned with the underlying point density map, even in regions with limited or missing wall evidence. Correspondingly, the reconstructed polygons evolve from coarse outlines to precise, watertight layouts. This illustrates the effectiveness of our progressive decoding strategy in improving both geometric accuracy and topological consistency.
B.2 More Reconstruction result on Structured3D
B.3 Failure cases on Structure3D
B.4 More Reconstruction result on SceneCAD
B.5 Extended Evaluation
Table 5: Sensitivity of reconstruction accuracy to threshold values. $\dagger$: the threshold we used in our experiments.
Eps Room Corner Angle
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
0.2 99.59 98.69 99.14 95.04 88.57 91.70 92.57 86.40 89.39
[…]
0.05 99.59 98.69 99.14 94.96 88.55 91.65 92.46 86.34 89.30
0.01 99.59 98.69 99.14 94.95 88.58 91.66 92.40 86.34 89.27
0.0001 99.59 98.69 99.14 94.94 88.60 91.67 92.38 86.34 89.26
Figure 13: Impact of the number of tokens on model complexity and efficiency. As the token count increases, the number of parameters, FLOPs, inference time, and GPU memory usage grow nearly quadratically, reflecting the scalability characteristics of transformer-based architectures.
Sensitivity to threshold in edge-to-polygon conversion.
The edge-to-polygon conversion’s sensitivity to the threshold (epsilon) is minimal as shown in Table 5, varying epsilon from 0.0001 to 0.2× edge length changes Room metrics by <0.01% and other metrics by less than 0.05%, confirming stability and negligible impact on accuracy. This threshold-based approach offers a practical trade-off between theoretical rigor and engineering feasibility.
Computational cost analysis.
We analyze the computational cost of our model with respect to the number of tokens used in the decoder. The metrics include the number of parameters, floating-point operations (FLOPs), inference time per image, and GPU memory usage for a batch of 10. Figure 13 summarizes these metrics for different token counts. Increasing the number of tokens leads to a higher number of parameters and FLOPs, longer inference time, and increased GPU memory consumption. Our method is inherently scalable and built on a general framework. The primary constraint is token length, which is bounded by GPU memory. Scaling is feasible by increasing the token count, though it come with quadratic complexity (O(n²) in both runtime and memory), as is typical for transformer-based models.